Prove that using proof by contradiction.

√2 + √6 < √15


1
Expert's answer
2021-10-25T16:27:14-0400

Assume the result is False; i.e. assume that "\sqrt{6}+\sqrt{2}\geq\sqrt{15}."

Then


"\sqrt{6}+\sqrt{2}\geq\sqrt{15}>0=>(\sqrt{6}+\sqrt{2})^2\geq(\sqrt{15})^2"

"=>6+2\sqrt{12}+2\geq15=>4\sqrt{3}\geq7"

"=>(4\sqrt{3})^2\geq(7)^2=>48\geq49,"

which is contradiction.

Hence we have proved that "\sqrt{6}+\sqrt{2}<\sqrt{15}."